# On putative self-similarity for incompressible 3D Euler

### v1

@@ 1. Introduction
Beyond a general intrinsic interest in lower bounds for $\gamma$ as a benchmark for ongoing computational and analytical studies, in the second half of the paper we focus on the hypothetical case when $\gamma$ lies below the parabolic threshold of $1/2$, relevant for incompressible 3D Navier-Stokes. At $\gamma=1/2$, the advection operator $\partial_{t}+u\cdot\nabla$ and the dissipative operator $\Delta$ are in exact balance when acting on vorticities in the form (1.2). Liouville theorems  [34, 44, 40, 8] rule out certain globally self-similar solutions for the 3D incompressible Navier-Stokes equations with $\gamma=1/2$. Under reasonable assumptions, $\gamma>\frac{1}{2}$ cannot be the exponent of an approximate self-similar blow up for 3D incompressible Navier-Stokes equations (see Proposition 3.1). For self-similar scaling with $\gamma<1/2$, the viscous term yields a vanishingly small force competing with the Euler nonlinearity. If self-similar solutions of Euler equations with $\gamma<\frac{1}{2}$ are found, they can be used to obtain 3D Navier-Stokes blow up as a perturbation of the 3D Euler one, allowing the inviscid blowup to be used to find a viscous one. This is indeed a rigorously proven fact in the realm of compressible flows: globally self-similar implosion singularities for 3D compressible isentropic Euler equations can be used to provide asymptotically self-similar implosion singularities for 3D compressible barotropic Navier-Stokes equations with constant viscosity coefficients; see [35, 36] and [3, 41]. In the process of establishing the blow up  [36, 3, 41], in addition to having $\gamma<1/2$ the authors prove and use that the inviscid similarity profile has infinite regularity (real-analyticity) in a neighborhood of all sonic points (stagnation points on the self-similar fast-acoustic characteristics).
@@
The paper [25] identifies an outgoing property (cf. (3.34) with $c_{*}>0$) as being an essential ingredient for the existence of self-similar profiles. This outgoing property quantifies the statement that all self-similar Lagrangian trajectories that originate from nonzero labels must escape to infinity as (self-similar) time diverges. In particular, this condition implies that the self-similar Lagrangian flow has no stagnation points, except for the trivial one the origin.
In this paper we prove that if an outgoing globally self-similar smooth solution to incompressible 3D Euler equations exists, then we must have $\gamma\geq 1/2$; see Theorem 3.7, Theorem 3.9, and Theorem 4.3, and Theorem 4.6. Thus, in this case, the approach proposed to produce incompressible Navier-Stokes singularities via Euler self-similar solutions must fail.
When restricting our analysis to axisymmetric similarity profiles, we prove stronger results. If the velocity has nonzero swirl at a stagnation point of the self-similar Lagrangian flow map, then $\gamma=1/2$. Alternatively, irrespective of the swirl value, if the velocity field is sufficiently smooth in a neighborhood of all stagnation points of the self-similar Lagrangian flow map, and if all these stagnation points lie on the axis of symmetry, then $\gamma\geq 1/2$; see Theorem 4.4.
@@
Our proofs are based on basic, well known properties of incompressible Euler and Navier-Stokes equations. We chose to present rigorous results, with explicit minimal assumptions. The fact that $\gamma\geq\frac{2}{5}$ for finite kinetic energy flows, is such a result. The relevance of the exponent $\gamma=\frac{1}{2}$ in Euler flows is one of the main messages of this paper. This exponent, which is natural for Navier-Stokes self-similarity, is natural for Euler equations as well. This is due to the invariance of circulation. a quantity with dimension of kinematic viscosity.
@@ 2. On asymptotically self-similar blowup with finite kinetic energy
In particular, if $L(t)=L_{0}(1-\frac{t}{T_{*}})^{\gamma}$ as in (1.2), then $\gamma>\frac{2}{5}$ is necessary for blow up. Here we revisit this result, without assuming the existence of a globally self-similar profile; instead, we assume only local behavior consistent with self-similarity. Furthermore, we link the putative exponent $\gamma$ to the behavior of the velocity away from the local self-similar behavior. In [7] it was shown that self-similar blow up cannot occur if the vorticity profile decays at infinity fast enough. This is not what we do here, our assumed behavior does not invoke a profile, but if a profile existed, it would have required only that the profile’s first derivatives be bounded in a fixed ball. We make assumptions about the nature of the true solution of (1.1), which would be automatically satisfied if a self-similar $C^{1}$-smooth vorticity profile existed.
@@
We consider a smooth cutoff function $\chi\colon\mathbb{R}_{+}\to\mathbb{R}$, with $\chi\equiv 1$ on $[0,1]$, and $\chi\equiv 0$ on $2,\infty]$. Let $R(t)>0$, to be determined later. We decompose $\alpha=\alpha_{\rm in}+\alpha_{\rm out}$, where
@@
$:=\frac{3}{4\pi}\int_{\mathbb{R}^{3}}\left(\hat{y}\cdot\xi(x,t)\right)\Bigl(\hat{y}\cdot\bigl((\omega(x+y,t)-\omega(x,t))\times\xi(x,t)\bigr)\Bigr)\chi\Bigl(\frac{|y|}{R(t)}\Bigr)\frac{dy}{|y|^{3}},$ (2.5)
@@ 3. On globally self-similar blowup
In order to obtain better lower bounds on the similarity exponent $\gamma$, we analyze the hypothetical case in which the 3D incompressible Euler equations (1.1), written in velocity form as
@@ 3.1. The self-similar ansatz
This ansatz is less restrictive than it seems. A more general form, $\omega(x,t)=A(t)\widetilde{\Omega}\left(\frac{x-x(t)}{\ell(t)},\tau(t)\right)$ leads to the form above. Indeed, writing the Euler equations in vorticity form and using the fact that the velocity and vorticity are related by the linear relation $\omega=\nabla\times u$, we see that the nonlinear term $u\cdot\nabla\omega-\omega\cdot\nabla u$ scales with amplitude $A^{2}$ under the general ansatz above. We thus arrive at the the ODE $\dot{A}=A^{2}$ where $\dot{A}$ is a time derivative  [^1] This time derivative can be with respect to an arbitrary internal clock, that is: $\frac{d}{dt}$ can be replaced by $\frac{1}{f(t)}\frac{d}{dt}$ where $f(t)$ is an arbitrary non-vanishing function.. This leads inevitably to $A(t)=(T-t)^{-1}$. The rest follows dividing the Euler equation by $A^{2}$ and taking the simplest situation, when the ensuing coefficients are constant. Then the constant coefficient in front of $y\cdot\nabla\Omega$ leads to $\ell(t)=\ell_{0}\left(1-\frac{t}{T}\right)^{\gamma}$. The term due to $x(t)$ is integrated, because $A$ and $\ell$ are known. It follows that $x(t)=x_{*}-\frac{c}{\gamma}\ell(t)$ where $c$ is a constant, and this is dealt with by translating $\widetilde{\Omega}$.
@@ 3.1.2. Space rescaling and normalization
Note that if $U$ and $\Omega$ are a solution of (3.4) on $\mathbb{R}^{3}$, then $U_{\lambda}(y)=\lambda U(\tfrac{y}{\lambda})$ and $\Omega_{\lambda}(y)=\Omega(\tfrac{y}{\lambda})$, are also a solution of (3.4) on $\mathbb{R}^{3}$. Under this rescaling, we have [^3] As we shall see later (cf. (3.8)), $\Omega$ is only expected to lie in $L^{p}(\mathbb{R}^{3})$ when $\gamma p>3$.
@@ 3.1.3. Behavior of the profiles at space infinity
Thus, the assumption (3.9) is satisfied if $\omega$ is a smooth vorticity matched to an inner self-similar blow up ansatz with asymptotic behavior at infinity of the type (3.8). The main point is that both the inner blow up profile and the matching vorticity magnitude majorant $z(x,t)=|x-x(t)|^{-\frac{1}{\gamma}}\mathbf{1}_{\{\ell(t)\leq|x-x(t)|\leq 1\}}$ have the property that they belong to $L^{4}(dt,(L^{2}({\mathbb{R}}^{3}))$ if $\gamma>\frac{1}{2}$, which is a class of non-blow up.
@@ 3.3. Self-similar vortex stretching
In view of (3.8) we expect $|\Omega|$ to be integrable at infinity only for $p\gamma>3$, which is why we stated (3.12) with general $p$, as opposed to the standard $p=2$ formulation.
@@
where $\hat{z}=z/|z|$. Let $\chi=\chi(|z|)$ be a smooth radially decreasing cutoff function, with $\chi\equiv 1$ in $B_{1}$ and $\chi\equiv 0$ in $B_{2}^{\complement}$. Let $L>0$ be a length scale to be optimized later. Then, from (3.13), and (3.8), we obtain that for $p\gamma>3$ we obtain
$\left|\mathsf{A}(y)-\frac{3}{4\pi}\int_{\mathbb{R}^{3}}\left(\hat{z}\cdot\Xi(y)\right)\left[\hat{z}\cdot\left(\Omega(y+z)\times\Xi(y)\right)\right]\left(1-\chi\left(\tfrac{|z|}{L}\right)\right)\frac{dz}{|z|^{3}}\right|$
$\leq\frac{3^{1/p}(p-1)^{(p-1)/p}}{(4\pi)^{1/p}}\|\Omega\|_{L^{p}(\mathbb{R}^{3})}L^{-\frac{3}{p}}.$ (3.14)
@@
$|\mathsf{A}(y)|\leq\mathsf{C}_{p}\|\nabla\Omega\|_{L^{\infty}(\mathbb{R}^{3})}^{\frac{3}{p+3}}\|\Omega\|_{L^{p}(\mathbb{R}^{3})}^{\frac{p}{p+3}},\qquad\mathsf{C}_{p}:=6^{\frac{3}{p+3}}\bigl(\tfrac{3}{4\pi}\bigr)^{\frac{1}{p+3}}(p-1)^{\frac{p-1}{p+3}},$ (3.16)
@@
for any $p>\frac{3}{\gamma}$, where the constant $\mathsf{C}_{p}>0$ is as defined in (3.16).
@@
$\|\Omega\|_{L^{p}(\mathbb{R}^{3})}\geq\bigl(\tfrac{\pi}{162}\bigr)^{\frac{1}{p}}(p-1)^{-1+\frac{1}{p}},$ (3.18)
whenever $p\gamma>3$.
@@ 3.4. Self-similar Lagrangian particle trajectories
We note that for $a=0$ we have $Y(0,\tau)=0$ for all $\tau\in\mathbb{R}$; the origin is fixed point of the dynamical system induced by (3.20).
@@ 3.4.1. The self-similar Cauchy formula and the Weber formula
$e^{(1-2\gamma)\tau}((\nabla Y)^{T}U(Y))(a,t)-U(a)=-\nabla\int_{0}^{\tau}e^{(1-2\gamma)s}\left(P(Y)-\frac{|U(Y)|^{2}}{2}\right)ds$
@@ 3.4.2. The self-similar Kelvin circulation theorem
$\begin{array}[]{l}e^{(1-2\gamma)\tau}\Gamma_{\rm{ss}}(\tau)=e^{(1-2\gamma)\tau}\oint_{C(\tau)}U(y)dy\\ =e^{(1-2\gamma)\tau}\int_{\mathbb{T}}U^{j}(Y(a(\lambda),\tau)\;(\partial_{i}Y^{j})(a(\lambda),\tau)\dot{a}^{i}(\lambda)d\lambda\\ =\Gamma_{\rm{ss}}(0)\end{array}$ (3.28)
@@ 3.4.3. The self-similar Bernoulli function
$\frac{d}{d\tau}\mathcal{H}(Y(a,\tau))=(2\gamma-1)|V(Y(a,\tau))|^{2}.$
@@ 3.5. An outgoing property?
We note that all trajectories with sufficiently large labels escape to infinity. Indeed, the bound (3.8) implies that for all $y\in\mathbb{R}^{3}$ such that $|y|\geq\bigl(\tfrac{2\mathsf{C}_{\flat}}{\gamma}\bigr)^{\gamma}$ we have $|U(y)\cdot y|\leq\frac{\gamma}{2}|y|^{2}$, and therefore $(\gamma y+U(y))\cdot y\geq\frac{\gamma}{2}|y|^{2}$. From (3.20) we then immediately obtain that if $|a|\geq(\frac{2\mathsf{C}_{\flat}}{\gamma})^{\gamma}$, then $|Y(a,\tau)|$ is a strictly increasing function of $\tau$, and we have the bounds $|a|e^{\frac{\gamma}{2}\tau}\leq|Y(a,\tau)|\leq|a|e^{\frac{3\gamma}{2}\tau}$, and $e^{\frac{\gamma}{2}\tau}\leq|\nabla_{a}Y(a,\tau)|\leq e^{\frac{3\gamma}{2}\tau}$, for all $\tau\geq 0$.
The exponential escape to infinity of self-similar Lagrangian flow maps (as exhibited above) is well-known to play a fundamental role, both in proving both the existence and the stability of self-similar profiles. In the context of the compressible Euler equations, see [5, 35, 3, 10]. In our incompressible 3D Euler setting we only know that trajectories $Y(a,\tau)$ for which $|a|$ is sufficiently large escape to infinity, and that the trajectory emanating from $a=0$ is frozen at the origin.
A priori, we cannot rule out the existence of points $y\in\mathbb{R}^{3}\setminus\{0\}$ for which $\gamma y+U(y)=0$. In the compressible setting, such points exist—sonic points—and they are well known to cause tremendous headaches, such as severe instabilities, even within the class of radially symmetric solutions [29, 35, 3]. For incompressible 3D Euler equations, a global outgoing property, quantified as a lower bound of the type
$(\gamma y+U(y))\cdot y\geq c_{*}|y|^{2},\qquad\mbox{for all }y\in\mathbb{R}^{3},\qquad\mbox{for some }c_{*}\geq 0,$ (3.34)
has been recently identified by Elgindi [25] as one of the two fundamental assumptions needed to prove the existence of solutions to (3.4); see [25, Hypothesis 6.7], where (3.34) is stated with $c_{*}>0$. In this section we allow the degenerate case $c_{*}=0$ to occur. This degeneracy can occur either at $y=0$ (when $\gamma|y|^{2}+y\cdot U(y)$ vanishes as $|y|^{2n}$ for some $n\geq 2$ as $|y|\to 0$), or the degeneracy can occur at some $y\neq 0$; in the later case, we assume that the zeros of $\gamma y+U(y)$ are isolated.
@@
Let $\Omega$ be a $C^{1}$ smooth globally self-similar vorticity profile for 3D incompressible Euler. If $\Omega(0)\neq 0$, and the outgoing property (3.34) holds in an open neighborhood of $y=0$ for some $c_{*}\geq 0$, then
@@
$\Omega+\bigl((\gamma y+U)\cdot\nabla\bigr)\Omega=\mathbb{S}\Omega,$ (3.35)
where $\mathbb{S}=\frac{1}{2}((\nabla U)+(\nabla U)^{\intercal})$ is the rate of strain matrix. If $y_{*}\in\mathbb{R}^{3}$ is a zero of $\gamma y+U(y)$, restricting (3.35) to $y=y_{*}$, if $\Omega(y_{*})\neq 0$, we deduce that the matrix

### v2

@@ 1. Introduction
Beyond a general intrinsic interest in lower bounds for $\gamma$ as a benchmark for ongoing computational and analytical studies, in the second half of the paper we focus on the hypothetical case when $\gamma$ lies below the parabolic threshold of $1/2$, relevant for incompressible 3D Navier-Stokes. At $\gamma=1/2$, the advection operator $\partial_{t}+u\cdot\nabla$ and the dissipative operator $\Delta$ are in exact balance when acting on vorticities in the form (1.2). Liouville theorems  [36, 44, 40, 8] rule out certain globally self-similar solutions for the 3D incompressible Navier-Stokes equations with $\gamma=1/2$. Under reasonable assumptions, $\gamma>1/2$ cannot be the exponent of an approximate self-similar blow up for 3D incompressible Navier-Stokes equations (see Proposition 3.1). For self-similar scaling with $\gamma<1/2$, the viscous term yields a vanishingly small force competing with the Euler nonlinearity. If self-similar solutions of Euler equations with $\gamma<1/2$ are found, they can be used to obtain 3D Navier-Stokes blow up as a perturbation of the 3D Euler one, allowing the inviscid blowup to be used to find a viscous one. This is indeed a rigorously proven fact in the realm of compressible flows: globally self-similar implosion singularities for 3D compressible isentropic Euler equations can be used to provide asymptotically self-similar implosion singularities for 3D compressible barotropic Navier-Stokes equations with constant viscosity coefficients; see [34, 35] and [3, 41]. In the process of establishing the blow up  [35, 3, 41], in addition to having $\gamma<1/2$ the authors prove and use that the inviscid similarity profile has infinite regularity (real-analyticity) in a neighborhood of all sonic points (stagnation points on the self-similar fast-acoustic characteristics).
@@
The paper [25] identifies an outgoing property (cf. (3.36)) as being an essential ingredient for the existence of self-similar profiles. This outgoing property quantifies the statement that all self-similar Lagrangian trajectories that originate from nonzero labels must escape to infinity as (self-similar) time diverges. In particular, this condition implies that the self-similar Lagrangian flow has no stagnation points, except for the trivial one at the origin.
In this paper we prove that if an outgoing globally self-similar smooth solution to incompressible 3D Euler equations exists, then we must have $\gamma\geq 1/2$; see Theorem 3.8, Theorem 3.10, and Theorem 4.3. Thus, in this case, the approach proposed to produce incompressible Navier-Stokes singularities via Euler self-similar solutions must fail.
When restricting our analysis to axisymmetric similarity profiles, we prove stronger results. If the velocity has nonzero swirl at a stagnation point of the self-similar Lagrangian flow map, then $\gamma=1/2$; see Theorem 4.4. Alternatively, irrespective of the swirl value, if the velocity field is $C^{1}$ smooth in a neighborhood of all stagnation points of the self-similar Lagrangian flow map, and if all these stagnation points lie on the axis of symmetry, then $\gamma\geq 1/2$; see Theorem 4.6.
@@
Our proofs are based on basic, well known properties of incompressible Euler and Navier-Stokes equations. We chose to present rigorous results, with explicit minimal assumptions. The fact that $\gamma\geq 2/5$ for finite kinetic energy flows, is such a result. The relevance of the exponent $\gamma=1/2$ in Euler flows is one of the main messages of this paper. This exponent, which is natural for Navier-Stokes self-similarity, is natural for Euler equations as well. This is due to the invariance of circulation, a quantity with dimension of kinematic viscosity.
@@ 2. On asymptotically self-similar blowup with finite kinetic energy
In particular, if $L(t)=L_{0}(1-\frac{t}{T_{*}})^{\gamma}$ as in (1.2), then $\gamma\geq 2/5$ is necessary for blow up. Here we revisit this result, without assuming the existence of a globally self-similar profile; instead, we assume only local behavior consistent with self-similarity. Furthermore, we link the putative exponent $\gamma$ to the behavior of the velocity away from the local self-similar behavior. In [7] it was shown that self-similar blow up cannot occur if the vorticity profile decays at infinity fast enough. This is not what we do here, our assumed behavior does not invoke a profile, but if a profile existed, it would have required only that the profile’s first derivatives be bounded in a fixed ball. We make assumptions about the nature of the true solution of (1.1), which would be automatically satisfied if a self-similar $C^{1}$-smooth vorticity profile existed.
@@
We consider a smooth cutoff function $\chi\colon\mathbb{R}_{+}\to\mathbb{R}$, with $\chi\equiv 1$ on $[0,1]$, and $\chi\equiv 0$ on $2,\infty)$. Let $R(t)>0$, to be determined later. We decompose $\alpha=\alpha_{\rm in}+\alpha_{\rm out}$, where
@@
$=\frac{3}{4\pi}\int_{\mathbb{R}^{3}}\left(\hat{y}\cdot\xi(x,t)\right)\Bigl(\hat{y}\cdot\bigl((\omega(x+y,t)-\omega(x,t))\times\xi(x,t)\bigr)\Bigr)\chi\Bigl(\frac{|y|}{R(t)}\Bigr)\frac{dy}{|y|^{3}},$ (2.5)
@@ 3. On globally self-similar blowup
In order to obtain better lower bounds on the similarity exponent $\gamma\geq 2/5$, we analyze the hypothetical case in which the 3D incompressible Euler equations (1.1), written in velocity form as
@@ 3.1. The self-similar ansatz
This ansatz is less restrictive than it seems. A more general form, $\omega(x,t)=A(t)\widetilde{\Omega}\left(\frac{x-x(t)}{\ell(t)},\tau(t)\right)$ leads to the form above. Indeed, writing the Euler equations in vorticity form and using the fact that the velocity and vorticity are related by the linear relation $\omega=\nabla\times u$, we see that the nonlinear term $u\cdot\nabla\omega-\omega\cdot\nabla u$ scales with amplitude $A^{2}$ under the general ansatz above. We thus arrive at the ODE $\dot{A}=A^{2}$ where $\dot{A}$ is a time derivative  [^1] This time derivative can be with respect to an arbitrary internal clock, that is: $\frac{d}{dt}$ can be replaced by $\frac{1}{f(t)}\frac{d}{dt}$ where $f(t)$ is an arbitrary non-vanishing function.. This leads inevitably to $A(t)=(T-t)^{-1}$. The rest follows dividing the Euler equation by $A^{2}$ and taking the simplest situation, when the ensuing coefficients are constant. Then the constant coefficient in front of $y\cdot\nabla\Omega$ leads to $\ell(t)=\ell_{0}\left(1-\frac{t}{T}\right)^{\gamma}$. The term due to $x(t)$ is integrated, because $A$ and $\ell$ are known. It follows that $x(t)=x_{*}-\frac{c}{\gamma}\ell(t)$ where $c$ is a constant, and this is dealt with by translating $\widetilde{\Omega}$.
@@ 3.1.2. Space rescaling and normalization
Note that if $U$ and $\Omega$ are a solution of (3.4) on $\mathbb{R}^{3}$, then $U_{\lambda}(y)=\lambda U(\tfrac{y}{\lambda})$ and $\Omega_{\lambda}(y)=\Omega(\tfrac{y}{\lambda})$, are also a solution of (3.4) on $\mathbb{R}^{3}$. Under this rescaling, we have [^3] As we shall see later (cf. (3.8)), $\Omega$ is only expected to lie in $L^{p}(\mathbb{R}^{3})$ when $p>3\gamma$.
@@ 3.1.3. Behavior of the profiles at space infinity
Thus, the assumption (3.9) is satisfied if $\omega$ is a smooth vorticity matched to an inner self-similar blow up ansatz with asymptotic behavior at infinity of the type (3.8). The main point is that both the inner blow up profile and the matching vorticity magnitude majorant $z(x,t)=|x-x(t)|^{-\frac{1}{\gamma}}\mathbf{1}_{\{\ell(t)\leq|x-x(t)|\leq 1\}}$ have the property that they belong to $L^{4}(dt;L^{2}({\mathbb{R}}^{3}))$ if $\gamma>\frac{1}{2}$, which is a class of non-blow up.
@@ 3.3. Self-similar vortex stretching
In view of (3.8) we expect $|\Omega|$ to be integrable at infinity only for $p>3\gamma$, which is why we stated (3.12) with general $p$, as opposed to the standard $p=2$ formulation.
@@
where $\hat{z}=z/|z|$. Let $\chi=\chi(|z|)$ be a smooth radially decreasing cutoff function, with $\chi\equiv 1$ in $B_{1}$ and $\chi\equiv 0$ in $B_{2}^{\complement}$. Let $L>0$ be a length scale to be optimized later. Then, from (3.13), and (3.8), we obtain that for $p>3\gamma$:
$\left|\mathsf{A}(y)-\frac{3}{4\pi}\int_{\mathbb{R}^{3}}\left(\hat{z}\cdot\Xi(y)\right)\left[\hat{z}\cdot\left(\Omega(y+z)\times\Xi(y)\right)\right]\chi\left(\tfrac{|z|}{L}\right)\frac{dz}{|z|^{3}}\right|$
$\qquad\leq\frac{3^{1/p}(p-1)^{(p-1)/p}}{(4\pi)^{1/p}}\|\Omega\|_{L^{p}(\mathbb{R}^{3})}L^{-\frac{3}{p}}.$ (3.14)
@@
$|\mathsf{A}(y)|\leq\mathsf{C}_{p}\|\nabla\Omega\|_{L^{\infty}(\mathbb{R}^{3})}^{\frac{3}{p+3}}\|\Omega\|_{L^{p}(\mathbb{R}^{3})}^{\frac{p}{p+3}},\qquad\mathsf{C}_{p}:=2\cdot 6^{\frac{3}{p+3}}\bigl(\tfrac{3}{4\pi}\bigr)^{\frac{1}{p+3}}(p-1)^{\frac{p-1}{p+3}},$ (3.16)
@@
for any $p>3\gamma$, where the constant $\mathsf{C}_{p}>0$ is as defined in (3.16).
@@
$\|\Omega\|_{L^{p}(\mathbb{R}^{3})}\geq\tfrac{1}{2}\bigl(\tfrac{\pi}{1296}\bigr)^{\frac{1}{p}}(p-1)^{-1+\frac{1}{p}},$ (3.18)
whenever $p>3\gamma$.
@@ 3.4. Self-similar Lagrangian particle trajectories
We note that for $a=0$ we have $Y(0,\tau)=0$ for all $\tau\in\mathbb{R}$; the origin is a fixed point of the dynamical system induced by (3.20).
@@ 3.4.1. The self-similar Cauchy formula and the Weber formula
$e^{(1-2\gamma)\tau}((\nabla Y)^{T}U(Y))(a,\tau)-U(a)=-\nabla\int_{0}^{\tau}e^{(1-2\gamma)s}\left(P(Y)-\frac{|U(Y)|^{2}}{2}\right)ds$
@@ 3.4.2. The self-similar Kelvin circulation theorem
$\begin{array}[]{l}e^{(1-2\gamma)\tau}\Gamma_{\rm{ss}}(\tau)=e^{(1-2\gamma)\tau}\oint_{C(\tau)}U(y)dy\\ =e^{(1-2\gamma)\tau}\int_{\mathbb{T}}U^{j}(Y(a(\lambda),\tau))\;(\partial_{i}Y^{j})(a(\lambda),\tau)\dot{a}^{i}(\lambda)d\lambda\\ =\Gamma_{\rm{ss}}(0)\end{array}$ (3.28)
@@ 3.4.3. The self-similar Bernoulli function
$\frac{d}{d\tau}\mathcal{H}(Y(a,\tau))=(2\gamma-1)|V(Y(a,\tau))|^{2}.$ (3.32)
@@ 3.5. An outgoing property?
We note that all trajectories with sufficiently large labels escape to infinity. Indeed, the bound (3.8) implies that for all $y\in\mathbb{R}^{3}$ such that
$|y|\geq R_{\flat}:=\max\left\{1,\Bigl(\frac{2\mathsf{C}_{\flat}}{\gamma}\Bigr)^{\gamma}\right\},$ (3.35)
we have $|U(y)\cdot y|\leq\frac{\gamma}{2}|y|^{2}$, and therefore $(\gamma y+U(y))\cdot y\geq\frac{\gamma}{2}|y|^{2}$. From (3.20) we then immediately obtain that if $|a|\geq R_{\flat}$, then $|Y(a,\tau)|$ is a strictly increasing function of $\tau$, and we have the bounds $|a|e^{\frac{\gamma}{2}\tau}\leq|Y(a,\tau)|\leq|a|e^{\frac{3\gamma}{2}\tau}$, and $e^{\frac{\gamma}{2}\tau}\leq|\nabla_{a}Y(a,\tau)|\leq e^{\frac{3\gamma}{2}\tau}$, for all $\tau\geq 0$.
The exponential escape to infinity of self-similar Lagrangian flow maps (as exhibited above) is well-known to play a fundamental role, both in proving the existence and the stability of self-similar profiles. In the context of the compressible Euler equations, see [5, 34, 3, 10]. In our incompressible 3D Euler setting we only know that trajectories $Y(a,\tau)$ for which $|a|$ is sufficiently large escape to infinity, and that the trajectory emanating from $a=0$ is frozen at the origin.
A priori, we cannot rule out the existence of points $y\in\mathbb{R}^{3}\setminus\{0\}$ for which $V(y)=\gamma y+U(y)=0$. In the compressible setting, such points exist—sonic points—and they are well known to cause tremendous headaches, such as severe instabilities, even within the class of radially symmetric solutions [29, 34, 3]. For incompressible 3D Euler equations, a global outgoing property, quantified as a lower bound of the type
$V(y)\cdot y=(\gamma y+U(y))\cdot y\geq c_{*}|y|^{2},\qquad\mbox{for all }y\in\mathbb{R}^{3},\qquad\mbox{for some }c_{*}>0,$ (3.36)
has been recently identified by Elgindi [25] as one of the two fundamental assumptions needed to prove the existence of solutions to (3.4); see [25, Hypothesis 6.7]. Assumption (3.36) implies that the only zero of $V$ occurs at $y=0$ (since $c_{*}>0$). In this section we want to allow for a generalized outgoing property, which allows for the degenerate case $c_{*}=0$, allows the function $V$ to vanish away from the origin, and only makes assumptions that are local, near the nodal set of $V$.
Definition 3.7.
We say that the field $V(y)=\gamma y+U(y)$ satisfies the local outgoing property if the nodal set
$\mathcal{N}_{V}:=\{y_{*}\in\mathbb{R}^{3}\colon V(y_{*})=0\}$
is finite, [^6] As discussed below (3.35), the bound (3.8) implies $\mathcal{N}_{V}\subset B_{R_{\flat}}(0)$. Hence, if we assume that $\mathcal{N}_{V}$ has no accumulation points, then this set must automatically be finite. and if there exists $c_{*}\geq 0$ and $\varepsilon_{*}>0$ such that for all $y_{*}\in\mathcal{N}_{V}$ we have
$V(y)\cdot(y-y_{*})\geq c_{*}|y-y_{*}|^{2},\qquad\mbox{when}\qquad|y-y_{*}|\leq\varepsilon_{*}.$ (3.37)
A few remarks concerning Definition 3.7 are in order:
  • Note that $\mathcal{N}_{V}\neq\emptyset$ because $0\in\mathcal{N}_{V}$ always.
  • The global outgoing property (3.36) implies the local outgoing property of Definition 3.7. Indeed, upon evaluating (3.36) at any $y_{*}\in\mathcal{N}_{V}$ we deduce that $0=V(y_{*})\cdot y_{*}\geq c_{*}|y_{*}|^{2}$; this shows $y_{*}=0$ since $c_{*}>0$. Thus $\mathcal{N}_{V}=\{0\}$. Condition (3.37) thus holds for any $\varepsilon_{*}>0$.
  • When $\mathcal{N}_{V}=\{0\}$, (3.37) generalizes (3.36) not just in the local-in-$y$ nature of the inequality; when $c_{*}=0$ it also allows for $V(y)\cdot y=\gamma|y|^{2}+y\cdot U(y)$ to vanish as $|y|^{2n}$ for $n\geq 2$ as $|y|\to 0$.
@@
Let $U$ be a $C^{2}$ smooth globally self-similar velocity profile for 3D incompressible Euler. If there exists $y_{*}\in\mathcal{N}_{V}$ such that $\Omega(y_{*})\neq 0$, and if the local outgoing property (3.37) holds in an open neighborhood of $y=y_{*}$ for some $c_{*}\geq 0$, then
@@
$\Omega+V\cdot\nabla\Omega=\mathbb{S}\Omega,$ (3.38)
where $\mathbb{S}=\frac{1}{2}((\nabla U)+(\nabla U)^{\intercal})$ is the rate of strain matrix. If $y_{*}\in\mathbb{R}^{3}$ is a zero of $V$, restricting (3.38) to $y=y_{*}$, we deduce that the matrix